# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-Edge AI Applications

### v3

@@ Abstract.
[…]
conference: ; ; Now at Microsoft AI. Now at xAI.
@@ 1. Introduction
GPU communication has become a key area of optimization for high-performance AI applications. For instance, modern Large Language Models (LLMs) (Dubey et al., 2024; Brown et al., 2020) are often designed to run a single inference task on tens of GPUs (spanning a few physical nodes) to distribute computation and minimize end-to-end latency. Training tasks often scale much further up to tens of thousands of GPUs (Dubey et al., 2024). GPUs are connected and communicate with each other through either intra-node (PCIe, NVLink (NVIDIA, 2025d), xGMI (AMD, 2023), etc.) or inter-node (Ethernet or InfiniBand (IB) (inf, 2023)) links. Efficient communication is challenging due to the collective communication patterns of AI workloads that frequently distribute and gather data across all GPUs simultaneously. For example, communication patterns such as AllReduce, AllGather, AllToAll, Broadcast, and ReduceScatter (Wikipedia, 2025) are frequently used in modern AI applications to distribute computing workload over many GPUs and collect computed results into a single GPU.
A substantial portion of the end-to-end latency is often attributed to GPU communication in real workloads, usually in the 10% - 40% range of end-to-end LLM workloads. For example, a GPT-3 (Brown et al., 2020) model inference may spend 30% of the end-to-end time in AllReduce communication. Many efforts have been made to optimize GPU communication, such as developing efficient communication algorithms (routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025a; AMD, 2025).
Despite these efforts, achieving high communication performance is still challenging and time consuming in real applications. Specifically, practitioners still need to write custom communication code to achieve the best performance, often from scratch. For instance, TensorRT-LLM (NVIDIA, 2025e), a popular LLM framework featured by NVIDIA in recent years, implements custom AllReduce communication methods from scratch. This is notable given that there already exists a popular GPU communication library developed by NVIDIA for many years, the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025a). The custom communication of TensorRT-LLM outperforms NCCL in a wide range of LLM scenarios, especially when the data size is relatively small, while TensorRT-LLM still uses NCCL for larger data sizes. This raises a simple question: Why build a new ad hoc stack from scratch, instead of developing on top of an existing stack?
Surprisingly, such ad hoc software development has already become a widely common practice in cutting-edge AI applications. From our empirical studies, we find that the root cause lies in the fast evolving hardware. The enormous computing demand by modern AI applications is pushing the industry to aggressively upgrade chips and interconnects, resulting in powerful but immature hardware that is dramatically different from the previous generation. Cutting-edge AI applications are usually targeted for deployment over the latest hardware to obtain the best performance and efficiency. Thus, it is not surprising for existing general-purpose libraries (such as NCCL) to perform sub-optimally for those applications because it takes time for those libraries to optimize for diverse workload scenarios. Consequently, production implementations go with custom quick-and-dirty optimization specific to their own workloads and hardware environments. This incurs redundant development efforts across applications and writes a lot of non-portable code, which leads to further development effort when a new hardware becomes available.
In this paper, we discuss an alternative interface design for AI software libraries to support portability and custom optimizations by reducing redundant efforts without harming the flexibility of customization. The key idea is to separate high-level abstractions and optimizations from primitive hardware abstractions based on two observations. First, existing hardware abstractions like NCCL primitives try to hide hardware features to make it easier to program, but may hinder optimizations. Instead, a low-level, primitive interface that minimally abstracts the hardware would make it easier for: (1) application developers to optimize and fine-tune by enabling precise control of the hardware, and (2) library developers to support new hardware features by only providing a shallow layer of abstraction over it. Second, existing collective communication libraries (like NCCL) and DSLs (domain-specific languages; like MSCCLang (Cowan et al., 2023)) provide ease of use, but may lack workload-specific optimizations. Such higher level abstractions could provide portability and enable users to trade-off performance for programming effort.
Based on these insights, we propose MSCCL++ (Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus), a novel GPU communication stack designed for high-performance AI applications. Unlike existing libraries, MSCCL++ exposes the most fundamental building blocks of communication as a user interface, namely the primitive interface (MSCCL++ Primitive API). This interface is very close to GPU hardware as a shallow layer over low-level GPU instructions, thus enabling quick adoption of latest hardware features with low effort. The primitive interface abstracts straightforward concepts of communication such as put, signal, wait, and flush. This helps translate low-level hardware behaviors into higher-level insights (e.g., bandwidth and latency) for optimizing communication. The primitive interface is designed to be very flexible to enable various optimizations that have been difficult to implement with existing stacks. It provides zero-copy, one-sided, and asynchronous abstractions for efficient communication both within and across nodes.
While the MSCCL++ Primitive API offers fine-grained optimization for GPU experts, MSCCL++ also provides higher-level interfaces that are easier to program. Inspired by MSCCLang (Cowan et al., 2023), we separate the declaration of communication algorithms, done using the MSCCL++ DSL API, from the underlying implementation of the algorithm, done by the MSCCL++ DSL Executor using MSCCL++ primitives. The MSCCL++ DSL accelerates the development of application-specific communication algorithms in new hardware environments. Lastly, for users with the least hardware expertise, we re-implement the popular NCCL API as-is (MSCCL++ Collective API) using MSCCL++ primitives so that applications can adopt it without changing the code at all.
@@
In our evaluation, we compare MSCCL++ with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs. We show that collective communication kernels like AllReduce and AllGather written using the MSCCL++ DSL API are up to $5.4\times$ faster than the other libraries. We also re-implement some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their counterparts. These algorithm implementations are included in the MSCCL++ Collective library. These benefits translate to AI inference workloads, yielding up to $15\%$ speedup in decodes over NCCL for a popular LLM, Llama2-70b (Touvron et al., 2023).
@@ 2.1. Collective Communication for AI
Modern AI applications are usually based on large machine learning (ML) models that require training and serving on multiple accelerator devices (GPUs in this paper) in a collective manner. They distribute the model and data across multiple GPUs to reduce the computation and memory requirement per GPU. This distribution requires GPUs to share intermediate results at some points during the computation, which is done using collective communication operations. AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums the input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU. ReduceScatter and AllGather are often used separately depending on the ML model architectures.
2.2. Existing Hardware Abstractions
Existing collective communication libraries implement conventional networking abstractions such as send and recv inside GPU kernels. This is a straightforward approach to generalize arbitrary communication algorithms in a hardware-agnostic manner. However, it is not the best choice for performance because such software abstractions are not well aligned with the GPU hardware abstractions. Instead, this paper proposes an API design from the perspective of GPU programming, which makes it easier to efficiently utilize GPU hardware. In this section, we provide an overview of the NCCL architecture and discuss its inefficiencies. This is representative of existing collective communication libraries for GPUs, because other popular libraries, such as RCCL (AMD, 2025) and MSCCL (Cowan et al., 2022), are designed based on NCCL and share the same limitations.
2.2.1. NCCL Architecture
NCCL provides a C++ API to setup connections between multiple GPUs and to run communication operations over GPUs. All participating processes (that usually hold one GPU each) of a distributed ML application initialize a NCCL handle to build connections beforehand, and then use the handle to call NCCL kernels (which perform collective operations on GPUs) during the execution. ML frameworks, such as PyTorch (Paszke et al., 2019) and TensorFlow (Abadi et al., 2016), wrap the NCCL API in Python, which are called by an ML model.
1global ringRS(in, nelem, ring)
2 send = ring.sendbuff; recv = ring.recvbuff;
3 sz = ring.buffSz;
4 prim = Primitives<half>
5 (tid, ring.buffSz, ring.prev, ring.next);
6
7 for (off = 0; off < nelem; off += sz)
8 //step 0: push data to the next GPU
9 idx = off + ring.ranks[ring.ranks-1]*sz;
10 prims.copy(in+idx, send, sz);
11 prims.send();
12
13 //k-2 steps: reduce and copy to next GPU
14 for (j = 2; j < ring.ranks; ++j)
15 rankDest = ring.ranks[ring.ranks-j];
16 idx = off + rankDest * buffSz;
17 prims.recv();
18 prim.reduce(in+idx, recv, send, sz, "+");
19 prims.send();
20
21 //step k-1: write the result for this rank
22 idx = off + ring.rank * buffSz;
23 prims.recv();
24 prims.reduce(recv, in+idx, sz, "+");
Figure 1. Ring ReduceScatter (simplified) in NCCL.
NCCL Initialization. During initialization, NCCL obtains following information: (i) the rank of each process, (ii) number of processes, (iii) number of distributed nodes in the cluster, (iv) number of GPUs per node, (v) all links between GPUs in a node, including PCIe, NVLink, and xGMI, and (vi) all links between nodes, like IB and Ethernet. NCCL uses this information to create communication topologies such as ring and tree with bi-directional links. Then, NCCL allocates a send buffer that stores data to send to destination GPU and a receive buffer that stores data received from a source GPU. Instead, users can also register their own send and receive buffers (using ncclMemAlloc()). The metadata of the send and receive buffers on a GPU are shared with all peer GPUs, so that GPUs can either peer-to-peer access buffers of another GPU (through PCIe, NVLink, or xGMI) or use external hardware to copy data (through IB or Ethernet).
NCCL Kernel. NCCL implements a GPU kernel for all collective operations and communication algorithms (ring, tree, etc.). It chooses the best performing algorithm based on the data size of collective communication, namely the message size. NCCL kernels are built using four building block operations, called primitives, which are send, recv, copy, and reduce. (i) send copies data from the send buffer of source GPU to the receive buffer of destination GPU, (ii) recv waits until the transfer by source GPU’s send has finished execution and unblocks send of the source GPU, (iii) copy copies source to destination buffer on the same GPU using GPU threads, and (iv) reduce does elementwise operation, like addition, on two buffers on the same GPU and write the result to another buffer. Calls to send blocks until destination GPUs have called recv to ensure that it is safe to write to send buffer. Additionally, NCCL defines extra operations that execute multiple primitives in a single fused function. However, this does not improve over the limitations we discuss in Section 2.2.2.
As an example, Figure 1 describes the NCCL kernel of a ring-based ReduceScatter for half precision floating point (FP16) using the NCCL primitives. The kernel takes the input buffer, number of elements, a ring of GPUs, and writes output to the input buffer. The kernel obtains pointers to send and receive buffers, size of the buffer, and initializes a primitive object with information about both previous and next GPUs in the ring (lines 25). The main loop of the kernel performs ReduceScatter in a batch of send/receive buffer sizes (lines 724). The loop first copies each batch of input data and sends it to the next GPU (lines 911). The loop then goes through all other ranks in the ring and for each rank (i) computes the corresponding offset in the input buffer for the rank, (ii) receives the partially reduced data from the previous GPU in the receive buffer, (iii) reduces this received data with the input offset and stores it in the send buffer, and (iv) sends this data to the next GPU in the rank (lines 1419). Finally, the GPU stores the data for the offset on this GPU (line 24).
2.2.2. Limitations of NCCL Primitives
The send and recv primitives are not flexible enough to utilize a single-instruction multiple-thread (SIMT) architecture like GPUs efficiently. As GPU threads are executed in groups (warps in CUDA), [^3] Unlike CPU where each thread runs on a separate context and can be context-switched, GPU threads are grouped and executed altogether for a single instruction. Otherwise, it may underutilize the highly-parallel cores and the large memory bandwidth. GPU kernels usually optimize the workload case-by-case to carefully distribute tasks among threads and minimize synchronization overhead. Unfortunately, NCCL only implements a static abstraction that groups 128 - 640 threads to collectively call a single primitive and synchronize them at the end of the call. This static abstraction is not flexible enough to optimize the workload for different links and data sizes, and may introduce unnecessary overheads. We discuss the limitations of NCCL primitives below.
Wasted GPU cycles. The send and recv primitives block the calling threads until the data transfer is completed, and this may waste a bunch of GPU threads’ cycles in a busy-wait while loop. If the underlying link supports peer-to-peer memory access, send directly copies local data into the receive buffer of the destination GPU, which can efficiently utilize parallel GPU threads. However, if the link uses external hardware (e.g., NIC) for the transfer, send only wakes up a CPU thread (writes a few-byte flag) that initiates the transfer and busy-waits until the transfer is completed (see Figure 2). This is a light-weight task that only needs a single GPU thread, and it is a big waste of cycles to block many threads for this. An alternative would be making the send and recv asynchronous, which is also against the existing abstractions in NCCL.
Inflexible Synchronization. To ensure data consistency, the NCCL primitives are designed to be self-synchronous. Such synchronizations are often too conservative to realize correct communication semantics. Consider a loop where $N$ GPUs produce their own data, and each GPU consumes data from all other GPUs, and starts all over again. When a single buffer is written by producers and read by consumers, two types of cross-GPU barriers are required: after each producer to ensure the following remote consumer sees consistent data, and after each consumer to prevent the following remote producer from overwriting what is being read by this consumer. To reduce synchronization overhead, we can use two buffers in rotation and only keep the first type of barriers which is sufficient to prevent overwriting. However, this is not possible with self-synchronized NCCL primitives.
Figure 2. Workflow of the NCCL send and recv primitives between two GPUs connected using InfiniBand (IB). Even though IB NICs can directly read and write GPU memory, we need the CPU to initiate the transfer (ibv_post_send) and wait for the completion (ibv_poll_cq). GPU 0 waits before calling send for next batch of data, and GPU 1 waits before the receive buffer is ready.
Interconnect Optimizations. A link can allow different ways of transferring data. However, NCCL supports only one mode of data transfer for each link. For example, intra-node links, like PCIe and NVLink, have two modes of data transfer: (i) thread-copy, where multiple GPU threads read from source GPU memory and write to destination GPU memory, and (ii) DMA-copy, where the CPU initiates the DMA engine of the GPU to copy from source memory to destination memory using cudaMemcpyDeviceToDevice. Compared to DMA-copy, thread-copy achieves lower latency, but often utilizes less bandwidth. For example, in our AllGather experiments over 8 NVIDIA A100 80G GPUs, thread-copy achieves only up to 227 GB/s NVLink bandwidth, while DMA-copy achieves 263 GB/s (+15.8%). However, NCCL uses only thread-copy unless it is infeasible by hardware. [^4] NCCL uses DMA copy only when peer-to-peer access between GPUs is not supported by the hardware configuration. Moreover, using DMA-copy frees GPU threads to do other work.
2.3. Existing Communication Libraries
NCCL API is designed to provide end-to-end implementations of collective communication operations. This would be what most applications expect from the library, but this is fundamentally limited to support diverse performance-sensitive applications. Like other GPU workloads do, collective communication over GPUs should be differently optimized for different scenarios, depending on the input size, GPU architecture, networking topology, and the application characteristics (Shah et al., 2023; Hwang et al., 2023a). It is very difficult for the library developers to provide the best performance for all cases in a timely manner.
MSCCL (Cowan et al., 2022) tackles this issue by allowing developers to implement their own custom communication algorithms using a DSL (Cowan et al., 2023). However, MSCCL is still based on the NCCL primitives and does not provide enough flexibility to fully utilize the hardware resources, as described in Section 2.2.2.

### v4

# MSCCL++: Rethinking GPU Communication Abstractions for AI Inference

@@ Abstract.
AI applications increasingly run on fast-evolving, heterogeneous hardware to maximize performance, but general-purpose libraries lag in supporting these features. Performance-minded programmers often build custom communication stacks that are fast but error-prone and non-portable.
This paper introduces MSCCL++, a design methodology for developing high-performance, portable communication kernels. It provides (1) a low-level, performance-preserving primitive interface that exposes minimal hardware abstractions while hiding the complexities of synchronization and consistency, (2) a higher-level DSL for application developers to implement workload-specific communication algorithms, and (3) a library of efficient algorithms implementing the standard collective API, enabling adoption by users with minimal expertise.
Compared to state-of-the-art baselines, MSCCL++ achieves geomean speedups of $1.7\times$ (up to $5.4\times$) for collective communication and $1.2\times$ (up to $1.38\times$) for AI inference workloads. MSCCL++ is in production of multiple AI services provided by Microsoft Azure, and has also been adopted by RCCL, the GPU collective communication library maintained by AMD. MSCCL++ is open source and available at https://github.com/microsoft/mscclpp. Our two years of experience with MSCCL++ suggests that its abstractions are robust, enabling support for new hardware features, such as multimem, within weeks of development.
GPU Communication; Collective Communication; AI Inference; Multi-layered Programming Abstraction
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790188isbn: 979-8-4007-2359-9/2026/03ccs: Software and its engineering Domain specific languagesccs: Software and its engineering API languagesccs: Software and its engineering Requirements analysisccs: Networks Programming interfacesccs: Software and its engineering Abstraction, modeling and modularityccs: Software and its engineering Software performanceccs: Software and its engineering Software usability
@@ 1. Introduction
GPU communication has emerged as a critical bottleneck and an opportunity for optimization in AI systems. For instance, communication kernels account for approximately 10%-40% in many real-world Large Language Model (LLM) inference tasks (NVIDIA, 2025f; Gond et al., 2025; Hwang et al., 2023a). Recent research has addressed communication bottlenecks, including developing efficient communication algorithms (e.g., routing paths and transfer scheduling) (Shah et al., 2023; Cao et al., 2024; Liu et al., 2024), overlapping communication with computation (Jangda et al., 2022; Wang et al., 2023), and other lower-level optimizations in the stack (NVIDIA, 2025d; AMD, 2025).
Despite these efforts, achieving high communication performance is still challenging and time-consuming in real applications. A common practice today is for practitioners to write custom communication code from scratch to achieve maximum performance without using standard libraries such as the NVIDIA Collective Communication Library (NCCL) (NVIDIA, 2025d). For instance, NVIDIA’s TensorRT-LLM (NVIDIA, 2025g) implements custom AllReduce communication kernels rather than relying on or updating NCCL. There are multiple reasons for doing so. First, achieving peak performance requires a careful balance between latency and bandwidth, tailored to the specific message sizes and communication patterns of a workload — something a general-purpose library, designed to support many workloads, cannot always optimize (Cowan et al., 2023). For example, TensorRT-LLM uses a custom AllReduce that outperforms NCCL on small data sizes, while falling back to NCCL for larger ones. Second, communication collectives often involve computations (e.g., reductions), and their performance depends on exploiting parallelism in ways that interact well with the necessary computation, further motivating workload-specific customization. Third, the enormous computational demand of modern AI applications is driving rapid evolution in both chips and interconnects, and practitioners rush to exploit these new capabilities long before general-purpose libraries such as NCCL are fully adapted.
The practice of writing custom communication code comes at a huge engineering cost. Developing these libraries is challenging and error prone. New hardware capabilities are often immature or poorly documented, forcing developers to navigate low-level details with limited guidance. There are also complex synchronization and data consistency challenges that arise from interactions across GPUs, CPUs, and NICs, which must be carefully managed to avoid correctness issues and performance bottlenecks. Finally, achieving maximum efficiency typically demands the design of custom communication algorithms tailored to the specific workload and hardware, further increasing development complexity.
Our hypothesis to address these issues is that multi-layered programming abstractions are needed to provide performance, portability, and productivity at the same time. This paper proposes MSCCL++ [^1] MSCCL++stands for Microsoft Collective Communication Library ++, pronounced em-sickle plus-plus., a design methodology that evaluates this hypothesis for high-performance GPU communication. MSCCL++ provides portable, hierarchical abstractions that enable users to explore the trade-offs between productivity and performance. At its core, MSCCL++ provides a Primitive API that exposes minimal performance-preserving abstractions of hardware interconnect methods. These abstractions provide full flexibility for customization, including fine-grained overlapping of compute and communication, while hiding the complexities of synchronization and ensuring portability across hardware vendors and generations. Building on the Primitive API, MSCCL++ provides a high-level Domain-Specific Language (DSL) for specifying custom communication algorithms while the MSCCL++ DSL Executor handles implementation concerns and performs optimizations when executing the DSL program using the Primitive API. Using the DSL, MSCCL++ provides a library of collective algorithms optimized for different scenarios that together provide the same API as NCCL. Our design is thus based on a separation of concerns: (1) the Primitive API provides a portable substrate for expert programmers to exploit hardware-specific features for the best performance; (2) the DSL API offers a streamlined interface that allows application developers to implement workload-specific and network-topology-specific optimizations with minimal effort; and (3) the Collective API enables users to leverage these specialized implementations with no additional effort. These three layers offer abstractions that improve programmability while enabling progressive customization layer-by-layer to get the last ounce of performance.
Prior primitive interfaces not only are hardware-vendor-specific but also often mask certain hardware features, limiting opportunities for optimization; e.g., NCCL provides a low-level send-recv abstraction that prohibits developers from exploiting one-sided, asynchronous communication over interconnects, and NVSHMEM (NVIDIA, 2025e) provides a low-level shared-memory abstraction that prevents interconnect-specific optimizations such as using port-mapped DMA-copy on NVLink. In contrast, the MSCCL++ Primitive API exposes the fundamental building blocks of communication for different data transfer modes supported by hardware interconnects: memory-mapped I/O, port-mapped I/O, and switch-mapped I/O. While the building blocks are specific to the data transfer modes, they are portable across hardware vendors and generations supporting the corresponding data transfer mode. This interface supports one-sided and asynchronous abstractions for efficient communication both within and across nodes. The implementation of these primitives handles orchestration and synchronization across GPU, CPU, and NIC, providing clear data consistency and synchronization semantics.
The MSCCL++ DSL allows developers to specify communication algorithms while abstracting away implementation details. To improve programmability, it provides a thread-block-based API with a global view across all GPUs and exposes the underlying one-sided, asynchronous communication abstractions of the Primitive API. MSCCL++ automatically inserts intra-thread-block synchronizations, fuses instructions, and removes unnecessary memory accesses before lowering to an execution plan that is interpreted by the MSCCL++ DSL Executor. This design accelerates the development of application-specific communication algorithms for new hardware environments. Unlike MSCCLang (Cowan et al., 2023), our DSL retains the one-sided and asynchronous properties of the Primitive API and enables algorithms that overlap communication with computation. Using the DSL, we implement efficient communication algorithms, including those with computation-communication overlap, that are tuned for AI inference workloads. These algorithms are bundled into our library that provides the MSCCL++ Collective API.
@@
In our evaluation, we compare MSCCL++ with NCCL, ROCm Collective Communication Library (RCCL) (AMD, 2025) and Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) on NVIDIA A100, NVIDIA H100, and AMD MI300x GPUs. We show that collective communication kernels like AllReduce and AllGather implemented using the MSCCL++ DSL are on average $1.99\times$, $2.08\times$, and $1.43\times$ faster than NCCL, RCCL, and MSCCL respectively. We also develop some of these communication algorithms directly using the MSCCL++ Primitive API and show that those kernels are on average $3\%$ faster than their DSL counterparts. These benefits translate to AI inference workloads. LLM inference decodes are on average $1.11\times$ and $1.31\times$ faster by using MSCCL++ instead of NCCL in vLLM (Kwon et al., 2023) and SGLang (SGLang, 2025) respectively.
The MSCCL++ code base (Microsoft, 2025c) has been publicly available for two years. During this time, RCCL (AMD, 2025) has adopted both the MSCCL++ APIs and library as the default to support collective communication primitives on current and upcoming AMD hardware. SGLang (SGLang, 2025), a popular LLM inference framework, uses the MSCCL++ library for its collective communication. As a further validation of our design, when NVIDIA introduced support for multimem instructions for switch-mapped I/O (aggregation and multicast), incorporating this support in MSCCL++ (Section 4.3) required only $16$ person-weeks, allowing the rest of the MSCCL++ stack to leverage these primitives. Similarly, supporting NVIDIA’s multi-node NVLink (NVIDIA, 2025b) took only $2$ person-weeks.
@@ 2.1. Collective Communications
Communication in AI workloads often happens through collective operations in which multiple GPUs coordinate to transfer data among themselves and optionally performing computations on the data. AllReduce is a common collective communication operation that sums the partial results of all GPUs and broadcasts the computed result to all GPUs. AllReduce can be divided into two other collective operations: ReduceScatter and AllGather. ReduceScatter sums the input buffers on all GPUs and distributes the output buffer equally on all GPUs. AllGather collects a distributed buffer from all GPUs and stores the full buffer on each GPU.
2.2. Vendor Communication Libraries
GPU vendors provide communication libraries that implement the collective operations. NVIDIA Collective Communication Libraries (NCCL) (NVIDIA, 2025d) is an optimized library for collective communications on NVIDIA hardware. Similarly, AMD ROCm Collective Communication Library (RCCL) (AMD, 2025) is optimized for AMD hardware. Microsoft Collective Communication Library (MSCCL) (Cowan et al., 2022) enables executing custom communication algorithm optimized for a particular input size and topology of the hardware. We briefly discuss the design of NCCL (see Hu et al. (Hu et al., 2025) for a detailed description), since both RCCL and MSCCL are based on NCCL and they both have the same limitations as NCCL. [^2] RCCL has since adopted code and abstractions from MSCCL++.
NCCL implements a GPU kernel for all collective operations. The implementation chooses the communication algorithms (ring, tree, etc.) to implement the collective based on the data size. The core communication happens through four primitive operations: send, recv, copy, and reduce. NCCL allocates send and receive buffers to facilitate the transfer. These operations are synchronous. For instance, calls to send blocks until destination GPUs have called recv to ensure that it is safe to overwrite the send buffer.
2.3. Limitations of NCCL Primitives
NCCL primitives are synchronous, which improves programmability, but wastes GPU cycles in busy-wait loops and prevents overlapping of computation and communication. In addition, NCCL’s send and recv primitives work over internal send/receive buffers incurring local memory copies. Finally, by being a general-purpose library, NCCL hard codes many performance optimizations providing inflexibility to the users. For instance, NCCL provides limited capability to control the parallelism for performing reduction or copy operations during communication. Similarly, NCCL hardcodes a single transfer mode per link, even though interconnects like PCIe and NVLink support multiple modes of transfer with subtle performance tradeoffs.